OpenStudy (anonymous):

solve system of equations: L(W)=1200 pie(w/2pie)^2(L)=600

OpenStudy (anonymous):

i would love to but i cannot understand it

OpenStudy (anonymous):

\[L\times W=1200\] \[\pi\times (\frac{W}{\pi})^2\times L=600\]?

OpenStudy (anonymous):

yes, except its "w/2pie" squared. then i have to solve for L and W but i don't know how to do that using the substitution and elimination method